OpenStudy (anonymous):

A restaurant charges $12 for a buffet plus $2.60 for each drink. Another restaurant charges $15 for a buffet plus $2 for each drink. The system of equations is graphed below. Part 1: What is the solution? MY ANSWER: (5, 25) Part 2: Show your work to verify the solution for both equations. I DON'T KNOW THIS VERY WELL.

OpenStudy (anonymous):

Let C be cost, d be num of drinks C = 2.6d +12 C = 2d +15 solution is where they intersect or have equal costs 2.6d +12 = 2d +15 -2d -2d ----------------- 0.6d +12 = 15 -12 -12 ----------------- 0.6d = 3 --- -- 0.6 0.6 d = 5 So when they order 5 drinks, the cost is $25 at either restaurant

OpenStudy (anonymous):

Thanks :-)